# ASVD: Activation-aware Singular Value Decomposition for Compressing Large Language Models

### v3

@@ 3.4 KV Cache Compression
3.5 Sensitivity-based Truncation Rank Searching
In addressing the second challenge, the variations of singular values among different layers of LLMs, we introduce the Sensitivity-based Truncation Rank Searching (STRS) method. This challenge arises from the fact that different layers in LLMs exhibit varying degrees of sensitivity to information compression, which is reflected in the distribution of their singular values. STRS evaluates the layer sensitivity and decides the best truncation of singular values. In the realm of NLP, perplexity is a key metric for assessing how effectively a language model predicts a sequence of tokens [Brown et al., 2020]. Therefore, we use the reduction in perplexity on the calibration dataset to evaluate the sensitivity of each layer. Similar to post-training compression methods [Dettmers et al., 2022, Frantar et al., 2022, Frantar & Alistarh, 2023], we collect a small number of input token sequences as calibration dataset.
@@ 4 Experiments
4.1 Hyper-parameters Exploration
In our study, we initiate an exploration of hyper-parameters in ASVD, focusing on the activation channel significance metric and the control factor $\alpha$. This exploration is conducted on OPT-125m, a relatively small network that facilitates rapid evaluation. We rigorously explored the control factor $\alpha$, and two methods for quantifying activation significance: Absolute Mean Value of Input Activation and Absolute Maximum Value of Input Activation. More details can be found in the Appendix. In summary, based on our explorations, we chose $\alpha=0.5$ and the Absolute Mean Value method for setting the scaling matrix $S$ in the ASVD process in the following experiments.
@@ 4.2 Parameters Compression
We conducted a comprehensive evaluation of Adaptive Singular Value Decomposition (ASVD) on two LLaMA models: LLaMA-7b and LLaMA-2-7b [Touvron et al., 2023a, b], each comprising 7 billion parameters. For each model, we selected a calibration set with 32 samples, and each sample contains 2048 tokens, from the Wikitext dataset to assess the layer-wise sensitivity.
Figure 5: Perplexity trends following ASVD compression on LLaMA-2-13b, LLaMA-2-7b and LLaMA-7b.
@@
ASVD involves setting varying thresholds binary searching process, enabling us to observe the impact of different compression levels on model performance. This approach resulted in a range of compressed networks, each characterized by a unique compression ratio. We evaluated the performance of these compressed networks using perplexity as the primary metric, focusing on two datasets: Wikitext-2 and the Penn Treebank (PTB). The objective was to understand the relationship between compression ratio and model performance. The results, illustrated in Fig.5, reveal several key insights: (1) As the parameter ratio decreases, there is a corresponding increase in perplexity. As more parameters are compressed, the more error is introduced by SVD. This reduces representational capacity of the network, thereby resulting in higher perplexity scores. (2) A plateau region is observed when the parameter ratio exceeds 0.9. In this range, ASVD predominantly decompresses the less sensitive layers, resulting in minimal impact on prediction accuracy. (3) Below a parameter ratio of 0.85, there is a rapid increase in perplexity, indicating that the more sensitive layers are being decompressed to a lower truncation rank, adversely affecting the model’s performance.
We also present a detailed analysis of the performance of compressed networks at various parameter ratios. This analysis aims to elucidate the impact of different compression techniques on network efficiency and accuracy. Table 1 displays the performance metrics for two LLaMA models, LLaMA-7b and LLaMA-2-7b, under several compression scenarios. These metrics include Massive Multitask Language Understanding (MMLU) zero-shot evaluation, perplexity on the Wikitext dataset (wiki), and perplexity on the Penn Treebank dataset (ptb). Our observations indicate significant variations in performance as a function of the parameter ratio and the compression method employed. Specifically, the table highlights the performance of each model when using standard SVD, SVD with binary search for truncation ranks (SVD*), and ASVD at different parameter ratios ranging from 0.75 to 0.95.
@@ 4.3 KV Cache Compression
We evaluate the KV Cache compression by using ASVD to decompose k projection and v projection in transformer. Table 2 summarizes the results, showing the perplexities on the wikitext2 and Penn Treebank datasets for different KV cache compression ratios. It is evident that the perplexity values remain stable when the KV cache ratio is above 40%. When the ratio is lower than 40%, the performance of the network is decreased. These observations suggest that ASVD is effective to compress the KV cache without negatively impacting the model.
@@ 4.4 Integrating ASVD with Quantization
param ratio dataset FP16 INT8 INT6 NF4 FP16 INT8 INT6 NF4
1 wiki 5.47 5.48 5.52 5.65 4.88 4.88 4.90 4.98
ptb 20.82 20.82 20.97 22.70 29.15 29.12 29.03 30.31
0.95 wiki 5.64 5.64 5.70 5.83 4.94 4.95 4.97 5.08
ptb 23.98 23.95 25.10 35.91 31.93 31.67 31.97 33.89
0.9 wiki 5.93 5.94 6.02 6.20 5.12 5.11 5.16 5.31
ptb 32.63 32.19 32.83 40.82 34.03 33.64 33.66 34.93
0.85 wiki 6.74 6.73 6.90 7.43 5.54 5.56 5.62 5.90
ptb 59.84 63.76 66.88 427.59 39.32 40.02 38.15 44.49
This section explores the integration of ASVD with quantization techniques to compress large language models. We investigate the synergy between ASVD and simple quantization methods, Round-To-Nearest (RTN) and 4-bit NormalFloat (NF4) [Dettmers et al., 2023], as a preliminary step. Future work could extend this investigation to more advanced quantization methods. Our experimental framework involves two stages. Firstly, we apply ASVD to decompose the network. Subsequently, we quantize the decomposed weights. We employ per-channel asymmetric quantization, targeting two quantization levels: 8-bit and 6-bit. We employ NF4 for quantizing to 4-bit.
Table 3 summarizes the results of our experiments on LLaMA-7b and LLaMA-2-7b models. The following observations were made: 8-bit Quantization: The results indicate that 8-bit quantization has a negligible impact on model performance, both for the original and the ASVD-compressed networks. 6-bit Quantization: A slight degradation in prediction accuracy is observed with 6-bit quantization. This effect is more pronounced in networks with higher compression (lower parameter ratios), suggesting a compounded impact of compression and lower bit quantization. 4-bit Quantization: Upon quantizing the network into NF4, a further deterioration in prediction accuracy is observed. When param ratio is greater than 0.9, the performance decline attributed to quantization is approximately consistent with that of the non-decomposed network. For instance, in the case of LLaMA-2-7b without decomposition, the NF4 quantization results in a wiki PPL decrease of 0.18 compared to FP16. Similarly, at a param ratio of 0.95, the NF4 quantization yields a PPL decrease of 0.19 compared to FP16. In summary, the findings suggest that ASVD is compatible with weight quantization techniques.
@@ A.3 Inference Cost with Decomposed LLMs
Regarding the computational aspect, let’s consider the input matrix $\mathbf{X}\in\mathbb{R}^{n\times t}$ and the weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$. In the original linear layer, the matrix multiplication is represented as $\mathbf{Y}=\mathbf{X}\mathbf{W}$. The number of Multiply-Accumulate (MAC) operations, denoted as $C$, in the original linear layer can be computed as: $C=tmn$. After the ASVD decomposition, the matrix multiplication transforms into $\mathbf{Y}=\mathbf{A}_{k}\mathbf{B}_{k}$. To analyze the computational efficiency, we calculate the MAC operations, denoted as $C_{k}$, for this decomposed form. The computation for $C_{k}$ is given by: $C_{k}=tkm+tkn$
@@ 2  $p_{M}=\left\lfloor\frac{p_{L}+p_{H}}{2}\right\rfloor$
A.5 Difference with TensorGPT Xu et al. [2023].
@@ A.7 Hyper-parameters Exploration
From the data presented in the table, we observe that both activation-aware methods show superior performance compared to standard SVD+STRS. We also notice that Lower and higher values of $\alpha$ (0.1 and 2) exhibit lower performance, while mid-range values (0.5) lead to better performance, and the Absolute Mean Value method consistently outperforms the Absolute Max Value method. Therefore, based on our observations, we chose $\alpha=0.5$ and the Absolute Mean Value method for setting the scaling matrix $S$ in the ASVD process in the following experiments.
@@ A.8 Absorbing Singular Values
After we decompose a matrix via ASVD, we can represent the weight matrix as a product of three matrices, i.e., $\mathbf{W}\approx\mathbf{U}_{k}\mathbf{\Sigma}_{k}\mathbf{V}_{k}^{T}$. Thanks to the diagonal nature of matrix $\mathbf{\Sigma}_{k}$, we can further optimize the inference process. Specifically, we can efficiently absorb the singular values in $\mathbf{\Sigma}_{k}$ into the matrices $U_{k}$ and $V_{k}^{T}$. We achieve this fusion using the following strategy: $\mathbf{A}_{k}=\mathbf{U}_{k}\sqrt{\mathbf{\Sigma}_{k}}$ and $\mathbf{B}_{k}=\sqrt{\mathbf{\Sigma}_{k}}\mathbf{V}_{k}^{T}$. Consequently, we obtain a more computationally efficient matrix operation  [^1] See Appendix for details.:

### v4

@@ 4 Experiments
4.1 Settings
We conducted a comprehensive evaluation of Activation-aware Singular Value Decomposition (ASVD) on two series of Large Language Models (LLMs): LLaMA and LLaMA-2 [Touvron et al., 2023a; b]. Our experiments encompassed models ranging from 7 billion to 13 billion parameters. For each model, we selected a calibration set with 32 samples, and each sample contains 2048 tokens, from the Wikitext dataset to assess the layer-wise sensitivity. We explore two methods to set transform matrix $\mathbf{S}$. The first is the magnitude-based method (Eq.8), which is indicated by ASVD. We set $\alpha$ to 0.5 in our experiments  [^1] The exploration of hyper-parameter $\alpha$ can be found in the Appendix.. We also experimented with the Cholesky decomposition method (Eq.9) to set the transform matrix, denoted ASVD+ in our experiments.
@@ 4.2 Parameters Compression
Figure 5: Perplexity trends of ASVD compression on LLaMA-2-13b, LLaMA-2-7b and LLaMA-7b.
@@
Sensitivity-based Truncation Rank Searching (STRS in Sec.3.4) involves setting varying thresholds binary searching process, enabling us to observe the impact of different compression levels on model performance. This approach resulted in a range of compressed networks, each characterized by a unique compression ratio. We evaluated the performance of these compressed networks using perplexity as the primary metric, focusing on two datasets: Wikitext-2 (wiki) and the Penn Treebank (ptb). The results, illustrated in Fig.5, reveal several key insights: (1) As the parameter ratio decreases, there is a corresponding increase in perplexity. (2) A plateau region is observed when the parameter ratio exceeds 0.9. In this range, ASVD predominantly decompresses the less sensitive layers, resulting in minimal impact on prediction accuracy. (3) Below a parameter ratio 222parameter ratio 0.85 means compress the model size by 15%. of 0.85, there is a rapid increase in perplexity, indicating that the more sensitive layers are being decompressed to a lower truncation rank, adversely affecting the model’s performance.
We also present a detailed analysis of the performance of compressed networks at various parameter ratios. Table 1 displays the performance metrics for two LLaMA models, LLaMA-7b and LLaMA-2-7b, under several compression scenarios. These metrics include MMLU zero-shot evaluation, perplexity on the Wikitext dataset (wiki), and perplexity on the Penn Treebank dataset (ptb). Our observations reveal significant performance variations based on the parameter ratio and the compression method used. Specifically, the table highlights the performance of each model when using standard SVD, SVD with binary search for truncation ranks (SVD*), and ASVD at different parameter ratios ranging from 0.75 to 0.95.
We compare ASVD  and ASVD+ [^3] ASVD+ refer to ASVD with whitening method for obtaining the transformation matrix in Eq.9 with SVD-LLM [Wang et al., 2024]. The results in Table 2 show that ASVD+ can improve the performance of ASVD, especially when the compression ratio is high. ASVD+ also outperforms the SVD-LLM method, particularly when the compression ratio is less than 30%. This is because SVD-LLM does not consider the layer-wise differences. In contrast, our method uses Sensitivity-based Truncation Rank Searching to set each layer with a different compression ratio. However, when the compression ratio is larger than 30%, all of these methods can significantly improve the perplexity of the LLMs.
Table 2: The perplexity on wikitext2 of SVD-LLM, ASVD  and ASVD+. In this table, we take the setting of SVD-LLM that the lm head is not taken into consideration to compute the param ratio.
LLama-2-7b LLama-2-13b
Param ratio SVD-LLM ASVD ASVD+ Param ratio SVD-LLM ASVD ASVD+
0.95 6.93 5.64 5.56 0.95 5.70 4.94 4.93
0.9 7.27 5.93 5.74 0.9 5.94 5.12 5.03
0.85 7.76 6.74 6.10 0.85 6.24 5.54 5.26
0.8 8.38 8.91 6.86 0.8 6.66 6.53 5.77
0.75 9.30 18.97 8.38 0.75 7.22 8.71 6.54
0.7 10.67 159.21 10.62 0.7 8.00 20.82 7.82
0.65 12.82 1034.59 13.87 0.65 9.10 53.30 9.84
0.6 16.14 730.60 19.12 0.6 10.78 133.88 13.18
@@ 4.3 KV Cache Compression
We evaluate the KV Cache compression by using ASVD to decompose k projection and v projection in transformer [Vaswani et al., 2017]. Table 3 summarizes the results, showing the perplexities on the wikitext2 and Penn Treebank datasets for different KV cache compression ratios. It is evident that the perplexity values remain stable when the KV cache ratio is above 40%. When the ratio is lower than 40%, the performance of the network is decreased. These observations suggest that ASVD is effective to compress the KV cache without negatively impacting the model.
@@ 4.4 Integrating ASVD with Quantization
param ratio FP16
INT8
INT8
NF4
INT4
FP16
INT8
INT8
NF4
INT4
wiki
1 5.47 5.48 5.45 5.65 5.59 4.88 4.88 4.88 4.98 4.97
0.95 5.64 5.64 5.56 5.83 5.82 4.94 4.95 4.97 5.08 5.18
0.9 5.93 5.94 5.82 6.2 6.21 5.12 5.11 5.15 5.31 5.43
0.85 6.74 6.73 6.51 7.43 7.18 5.54 5.56 5.59 5.9 5.96
ptb
1 20.82 20.82 20.93 22.7 21.50 29.15 29.12 29.29 30.31 30.47
0.95 23.98 23.95 25.47 35.91 27.79 31.93 31.67 30.19 33.89 31.21
0.9 32.63 32.19 37.11 40.82 39.31 34.03 33.64 35.47 34.93 38.95
0.85 59.84 63.76 84.52 427.59 95.85 39.32 40.02 43.01 44.49 50.56
This section investigates the compatibility of ASVD with quantization techniques for compressing LLMs. We explore the integration of ASVD with different quantization methods. Simple quantization methods include Round-To-Nearest (RTN) and 4-bit NormalFloat (NF4) [Dettmers et al., 2023]. The advanced LLM quantization method is Activation-aware Weight Quantization (AWQ) [Lin et al., 2023]. Note that our study focuses on establishing the orthogonal property of ASVD to these basic quantization methods. Future work could extend this investigation to more advanced quantization techniques and other LLM compression approaches. Our experimental framework involves two stages. Firstly, we apply ASVD to decompose the network. Subsequently, we quantize the decomposed weights.
Table 4 summarizes the results of our experiments on LLaMA-2-7b, LLaMA-2-13b models [Touvron et al., 2023b]. The following observations were made: 8-bit Weight Quantization: The results indicate that 8-bit quantization has a negligible impact on model performance, both for the original and the ASVD-compressed networks. 4-bit weight Quantization: Upon quantizing the network into NF4 and INT4(AWQ), a further deterioration in prediction accuracy is observed. When param ratio is greater than 0.9, the performance decline attributed to quantization is approximately consistent with that of the non-decomposed network. We observe that the performance degradation of LLaMA-2-13b is less than that of LLaMA-2-7b, indicating that the larger model is more robust to compression. In summary, the findings suggest that ASVD is compatible with weight quantization techniques.
@@ 5 Conclusion
6 Reproducibility Statement
We have submitted the code for the experiments as part of the supplementary material. The code is anonymous and self-contained and includes detailed instructions to facilitate the replication of our experiments and findings. We also plan to publicly release the code, data, pretrained models, and any additional resources needed for the community to fully reproduce our work.
@@ A.1 Impact Statements and Limitations
Despite the remarkable achievements of the ASVD method in compressing large language models (LLMs), several limitations persist. One limitation arises when ASVD is combined with quantization techniques, which can lead to a decline in model performance. While 8-bit weight quantization has minimal effects on both original and ASVD-compressed networks, switching to 4-bit quantization can result in a slight decrease in predictive accuracy. Additionally, ASVD faces difficulties in compressing multi-layer perceptron (MLP) in LLMs, as these layers typically contain more parameters than self-attention mechanisms, resulting in increased computational burdens due to their high-dimensional feature mappings. Although ASVD effectively compresses the weights in multi-head attention (MHA) with fewer parameters, it struggles with MLP. Furthermore, the need to evaluate the sensitivity of each layer requires a forward propagation step to calculate perplexity, demanding significant computational resources.
@@ A.3 Inference Cost with Decomposed LLMs
Regarding the computational aspect, let’s consider the input matrix $\mathbf{X}\in\mathbb{R}^{n\times t}$ and the weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$. In the original linear layer, the matrix multiplication is represented as $\mathbf{Y}=\mathbf{W}\mathbf{X}$. The number of Multiply-Accumulate (MAC) operations, denoted as $C$, in the original linear layer can be computed as: $C=tmn$. After the ASVD decomposition, the matrix multiplication transforms into $\mathbf{Y}\approx\mathbf{U}_{k}^{\prime}\mathbf{\Sigma}_{k}^{\prime}\mathbf{V}_{k}^{\prime\prime}\mathbf{X}$. We can fuse the $\mathbf{\Sigma}_{k}$ into $\mathbf{U}_{k}^{\prime}$ and $\mathbf{V}_{k}^{\prime\prime}$. Then we have:
$\mathbf{Y}$ $\approx\mathbf{U}_{k}^{\prime}\mathbf{\Sigma}_{k}^{\prime}\mathbf{V}_{k}^{\prime\prime}\mathbf{X}$ (11)
$=(\mathbf{U}_{k}^{\prime}\sqrt{\mathbf{\Sigma}_{k}^{\prime}})(\sqrt{\mathbf{\Sigma}_{k}^{\prime}}\mathbf{V}_{k}^{\prime\prime})\mathbf{X}$ (12)
$=\mathbf{A}\mathbf{B}\mathbf{X}$ (13)
To analyze the computational efficiency, we calculate the MAC operations, denoted as $C_{k}$, for this decomposed form. The computation for $C_{k}$ is given by: $C_{k}=tkm+tkn$
@@ 2  $p_{M}=\left\lfloor\frac{p_{L}+p_{H}}{2}\right\rfloor$
A.5 Difference with TensorGPT.
@@ A.7 Hyper-parameters Exploration
From the data presented in the table, we observe that both activation-aware methods show superior performance compared to standard SVD+STRS. We also notice that Lower and higher values of $\alpha$ (0.1 and 2) exhibit lower performance, while mid-range values (0.5) lead to better performance, and the Absolute Mean Value method consistently outperforms the Absolute Max Value method. Therefore, based on our observations, we chose $\alpha=0.5$ and the Absolute Mean Value method for setting the transform matrix $\mathbf{S}$ in the ASVD process in the following experiments.
@@ A.8 Absorbing Singular Values
After we decompose a matrix via ASVD, we can represent the weight matrix as a product of three matrices, i.e., $\mathbf{W}\approx\mathbf{U}_{k}\mathbf{\Sigma}_{k}\mathbf{V}_{k}^{T}$. Thanks to the diagonal nature of matrix $\mathbf{\Sigma}_{k}$, we can further optimize the inference process. Specifically, we can efficiently absorb the singular values in $\mathbf{\Sigma}_{k}$ into the matrices $U_{k}$ and $V_{k}^{T}$. We achieve this fusion using the following strategy: $\mathbf{A}_{k}=\mathbf{U}_{k}\sqrt{\mathbf{\Sigma}_{k}}$ and $\mathbf{B}_{k}=\sqrt{\mathbf{\Sigma}_{k}}\mathbf{V}_{k}^{T}$. Consequently, we obtain a more computationally efficient matrix operation: